Commercial banks in United Economy have total deposits of AED 300 billion. Their reserves are AED 15 billion, two- thirds of which are with the Central Bank as deposits. There are AED 30 billion notes outside the banks. There are no coins!

Calculate-
a) The monetary base.
b) The banks’ reserve ratio.
c) Currency drain as a % of deposits.
1
Expert's answer
2014-12-02T13:21:29-0500
Deposits - 300 billion,
Total reserves - 15 billion,
Required reserves - 2/3*15 = 10 billion
Notes outside the banks - 30 billion
a) The monetary base is M = 300 + 15 + 30 = AED 345 billion
b) The banks’ reserve ratio rr = 10/300 = 0.033 or 3.3%
c) Currency drain as a % of deposits is 30/300 = 0.1 or 10%